OpenStudy (anonymous):

If alpha,beta are the roots of x^2-12x+7=0, find the values of: alpha^2+beta^2

zepdrix (zepdrix):

Hmm this does NOT factor, so let's try throwing it into the Quadratic Formula: \[\Large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (dape):

Well, if we already know the roots and call them \(\alpha\) and \(\beta\). We know they solve the equation, so we have \[\alpha^2=12\alpha-7 \\ \beta^2=12\beta-7\] Adding these two, we get \[ \alpha^2+\beta^2=12(\alpha+\beta)-14\]

OpenStudy (tkhunny):

I suppose, but this might be more interesting. \(\alpha\beta = 7\) \(\alpha + \beta = 12\) Squaring the second: \(\alpha^{2} + 2\alpha\beta + \beta^2 = 144\) Substitute from the first. Are you seeing it, yet? Interestingly, if you use the quadratic formula, square the two results, and add them, it results in \(\alpha^{2} + \beta^{2} = \dfrac{b^{2} - 2ac}{a^{2}} = ??\) Unique answers don't care how you find them!

OpenStudy (tkhunny):

Note: More interesting than Brute Force! About as interesting as @dape's solution.

OpenStudy (dape):

Right, if we use what @tkhunny said, namely \(\alpha+\beta=12\), we can put it into my formula and get \(\alpha^2+\beta^2=130\)

OpenStudy (dape):

But there are numerous ways of doing it as you can see when you take everything into account.

OpenStudy (anonymous):

yeah i see. thankyou everyone